# Resource-constrained Fairness

### v1

@@ 5 Discussion
Strategic adjustments may become necessary depending on the outcomes of this cost-benefit analysis. Up until now, we assumed the resource level to be fixed, but practitioners may find it feasible to increase or decrease this level, particularly if the resource levels were initially determined under an unfair model allocation. For instance, in healthcare, additional investments could be directed towards expanding screening facilities to counterbalance the decrease in true positives, effectively maintaining the detection rates prior to implementing fairness measures. Similarly, in educational settings such as student admission, institutions might adjust the size of admitted cohorts, either decreasing it to preserve the average quality, or increasing it to have the same number of graduates. Additionally, after doing this analysis, one could just select a resource level where the cost of fairness is low. Donahue and Kleinberg (2020) also discuss how increasing the level of available resources is a critical goal where advocacy and political action can play a key role. However, it will often not be possible to have a resource level equal to the whole population (Donahue and Kleinberg, 2020).
Another strategic approach could involve refining the precision for the protected groups by collecting more data or enhancing data quality, albeit at an initial cost. This investment might reduce the overall fairness cost, yielding a more profitable model over time. However, high quality data may simply not be available for all groups or infeasible to collect for a variety of legal, political, and practical reasons, meaning it cannot be assumed that collecting more or “better" data will necessarily solve inequalities in resource distribution (Perez, 2019; Pot et al., 2019; Wachter et al., 2020).
6 Conclusion and limitations
@@
In this work, we argued that fairness in machine learning should be investigated as a resource-constrained problem. We introduced the notion of resource-constrained fairness, and quantify the cost of fairness for a fixed level of resources (or a fixed selection rate). This allows us to investigate the actual cost of fairness, and not changes in performance metrics that are heavily influenced by a change in the overall selection rate. Our analysis closely aligns with the real-world challenges in enforcing fairness, where resources are often fixed, and allows organizations to have reasonable expectations about what costs to expect and why. Furthermore, we demonstrate that several dataset properties are relevant for the cost of fairness and that the cost of implementing fairness measures significantly depends on the decision-making context, particularly the used fairness metric and the level of available resources. This highlights the importance of evaluating the cost of fairness within the specific context of your problem.
However, we also see various limitations to our work. First, as already mentioned, assigning a cost to enforcing fairness is essentially a value judgement. We calculate the cost under the assumption that the ground truth is correct, while in reality, these labels can be biased as well. Despite this assumption, our results remain useful to understand potential trade-offs in enforcing fairness. Lastly, in our set-up we assume access to a static sensitive attribute. In reality, access to these attributes might be hard to obtain, as legal and ethical constraints may impose constraints on obtaining or utilizing certain sensitive information (Haeri and Zweig, 2020; Veale and Binns, 2017). In these scenarios, inferred characteristics could be used during deployment (Lohaus et al., 2022). Despite these limitations, this work still makes important contributions towards the field of Responsible AI in practice, by taking into account real-world constraints.
@@ Acknowledgments
Figure 6: Cost of fairness (loss in recall) for all datasets when enforcing demographic parity (DP) and equal opportunity (EO). We see that the cost of fairness for both metrics heavily depends on available resource level and thus the used selection rate.

### v3

@@ 4.4 Varying Global Selection Rates
Strategic adjustments in selection rate may be necessary depending on the outcomes of a cost-benefit analysis. Until now, we assumed the resource level to be fixed, but practitioners may find it feasible to increase or decrease this level, particularly if the resource levels were initially determined under an unfair model allocation. For instance, in healthcare, additional investments could be directed towards expanding screening facilities to counterbalance the decrease in true positives, effectively maintaining the detection rates prior to implementing fairness measures. Similarly, in educational settings such as student admission, institutions might adjust the size of admitted cohorts, either decreasing it to preserve the average quality, or increasing it to have the same number of graduates. Donahue and Kleinberg (2020) also discuss how increasing the level of available resources is a critical goal where advocacy and political action can play a key role. Another strategic approach could involve refining the precision for the protected groups by collecting more data or enhancing data quality, albeit at an initial cost. This investment might reduce the overall fairness cost, yielding a more profitable model over time. However, high quality data may simply not be available for all groups or infeasible to collect for a variety of legal, political, and practical reasons, meaning it cannot be assumed that collecting more or “better" data will necessarily solve inequalities in resource distribution (Perez, 2019; Pot et al., 2019; Wachter et al., 2020).
5 Conclusion
This work addresses the practical concern of enforcing fairness with a limited budget or, equivalently, as the selection rate/threshold varies. We demonstrated that standard methods for fairness do not enforce fairness for a varying threshold (Figure  2). We unified leveling up (Mittelstadt et al., 2023) and fairness under budgetary constraints, and have shown that the solutions for each approach coincide under constrained resources (Section 2.2). We have also shown how simple methods for setting per group thresholds can enforce fairness under these rate constraints (Section 3). Using this, we empirically measured the cost of fairness under rate constraints – in terms of decreased performance, or as an increase in global harm (Section 4); and derived theoretical bounds for these costs (Section 4.3). While the bounds are not intended to be tight, they are informative, and align with our empiric findings.
@@ 5 Conclusion
This allows us to investigate the actual cost of fairness, rather than changes in performance metrics that are heavily influenced by a change in the overall selection rate (Goethals et al., 2024). Compared to other fairness analyses, our analysis more closely aligns with real-world challenges, where resources are often fixed, and allows organizations to have reasonable expectations about what costs to expect and why. Furthermore, we demonstrated that the decision-making context matters, particularly the behavior of the classifier; the used fairness metric; and the level of available resources. This highlights the importance of evaluating the cost of fairness within the specific domain of use.
@@ Acknowledgments
Figure 5: Cost of fairness (loss in recall) for all datasets when enforcing demographic parity (DP) and equal opportunity (EO). We see that the cost of fairness for both metrics heavily depends on the available resource level and thus the used selection rate.